# Mitigating gain calibration errors from EoR observations with SKA1-Low AA*

### v1

@@ 3.3 Gaussian Process Regression (GPR)
where $\sigma^{2}$ is the variance, $\ell$ is the correlation lengthscale, and $\eta$ governs smoothness. Common choices include: RBF ($\eta\rightarrow\infty$), MATÉRN52 ($\eta=5/2$), MATÉRN32 ($\eta=3/2$), and exponential ($\eta=1/2$). For this work, we use RBF for sky foregrounds, Mat32/52 for mode-mixing, and exponential for the Hi signal.
@@ 3.3.1 Application
1% 10% 1% 10%
$\sigma_{\text{sky}}/\sigma_{n}$ $\infty$ $\mathbb{U}(0.1,5)$ $\mathbb{U}(0.1,1)$ 1.03 0.48
$l_{\text{sky}}$ (MHz) $\mathbb{U}(5,50)$ $\mathbb{U}(1,25)$ 24.11 3.03
$\sigma_{21}/\sigma_{n}$ 1/2 $\mathbb{U}(10^{-6},10^{-1})$ $\mathbb{U}(10^{-8},10^{-1})$ $10^{-3}$ $10^{-7}$
$l_{21}$ (MHz) $\Gamma(0.5,5)$ $\Gamma(0.5,5)$ 26.39 11.53
Table 2: The GPR kernel priors and posteriors for 1% and 10% gain calibration error cases.
Figure 6: Posterior distributions of GPR hyperparameters after MCMC sampling for 1% (left) and 10%(right) gain calibration errors. Contours represent confidence levels of 11.8%, 39.3%, 67.5%, and 87.4%.
@@ 4 Results
4.1 Case I: Gain Error of $\sim 0.1\%$
@@
In this case, we incorporate a gain calibration error of 0.1% and corrupt the simulated observed visibilities accordingly. The impact of a 0.1% gain calibration error on 1D PS is presented in Figure 4. In addition, the absolute fractional error of 2D PS, computed using Equation 2.7, is shown in the right panel of Figure 7. The 1D PS indicates that on larger scales ($k\leq 0.2$ Mpc-1), the residual power closely matches the target signal level, even without applying any mitigation. However, on smaller scales ($k>0.2$ Mpc-1), the residual power underestimates the true signal by nearly an order of magnitude. By applying foreground avoidance techniques using the horizon line, as shown in the left panel of Figure 7, the signal recovery improves significantly throughout the $0.05\leq k\leq 0.5$ Mpc-1 range, despite an estimated loss of $\sim$5% in SNR (see Section 3.1). For this specific scenario, neither subtraction nor hybrid mitigation methods were employed.
Figure 7: Power spectrum analysis after residual foreground mitigation for 0.1% gain calibration error. Left: Absolute fractional error map $\Delta P_{0.1\%}$ for sky-model subtracted visibilities, evaluated according to equation 2.7. The horizon limit (dotted black line) is used as the wedge boundary, with shaded regions indicating masked $k$-modes. Right: The square markers show the 1D PS before and after foreground avoidance. The grey-shaded region represents the true HI signal. Binning and error bars follow Figure 4.
4.2 Case II: Gain Error of $\sim 1\%$
@@
Figure 8: The 2D absolute fractional error maps $\Delta P_{i}(k_{\perp},k_{\parallel})$ for 1% (top row) and 10% (bottom row) gain calibration errors. Columns (from left to right) correspond to sky-model subtraction, PCA subtraction, and GPR subtraction. The dotted black lines indicate the wedge boundaries used for avoidance; the shaded areas represent the masked $k$-modes for that particular case. Bluer regions indicate better agreement with the true Hi signal. Corresponding 1D PS are shown in Figures 9 and 10.
@@
To further quantify the impact of calibration systematics, we introduce a gain calibration error of 1% and corrupt the simulated visibilities accordingly. The resulting cylindrical power spectrum is shown in the bottom-right panel of Figure 3, while the corresponding spherically averaged power spectrum is presented in Figure 4. A comparison between the residual power spectrum, affected by a 1% gain calibration error, and the true Hi power spectrum reveals substantial discrepancies across all $k$ modes, with residual power consistently exceeding the true signal by a significant margin. Relative to the case of 0.1% error, these deviations are substantially larger, clearly demonstrating that foreground avoidance alone is insufficient for reliable signal recovery under 1% gain calibration errors.
@@
The plots in the top row in Figure 8 show the variation in 2D absolute fractional error maps for 1% calibration gain error, after PCA subtraction, and after GPR subtraction, respectively. One can see that the subtraction techniques effectively reduce the absolute fractional error by up to an order of magnitude in the low $k_{\perp}$ range ($\leq 0.1$Mpc-1). Although GPR performs marginally better for higher $k_{\perp}$ modes than PCA, the mode-mixing component dominates, and we do not observe a significant improvement with either method. The corresponding spherically averaged power spectrum after the subtraction techniques is presented in the left plot of Figure 9.
Using avoidance around the horizon line or a 0.05 Mpc-1 buffer, our model fails to recover the signal on large scales. Avoidance with a buffer of 0.1 Mpc-1 (Figure 8, top-left) successfully recovers the signal but results in an average SNR loss of at least 30% (see Section 3.1), as indicated by the cyan markers in the right plot of Figure 9.
Lastly, using a hybrid technique of applying avoidance (with a 0.05 Mpc-1 buffer) to PCA/GPR residuals, it enables us to successfully recover the true signal Hiin the available $k$ modes, as represented by the green and blue markers in the right plot of Figure 9.
@@
Figure 9: Spherically averaged 1D PS after residual foreground mitigation for 1% gain calibration error. Left (Subtraction only): Star markers denote PCA and GPR-subtracted residuals. The shaded region represents the true Hi signal. Right (Hybrid): The cyan markers show avoidance using a 0.1 Mpc-1 buffer; green and blue markers represent hybrid (subtraction + 0.05 Mpc-1 buffer avoidance) cases. The orange and brown crossed-out regions represent the $k$-modes that are inaccessible due to avoidance with 0.05 and 0.1 Mpc-1 buffers, respectively. Markers for GPR are offset by 5% in $k$ for visibility.
4.3 Case III: Gain Error of $\sim 10\%$
@@
In this case, we corrupted the observed visibilities with a 10% gain calibration error to assess its impact on the recovered power spectrum. The effects are illustrated in the cylindrical power spectrum (bottom-left panel of Figure 3) and the spherically averaged power spectrum (Figure 4). We further examined the effectiveness of subtraction-based mitigation techniques in this high-error scenario. The plots in the bottom row in Figure 8 show the variation in the 2D absolute fractional error maps for 10% calibration gain error and after using subtraction methods. Applying GPR and PCA independently shows that subtraction alone is insufficient. GPR severely overestimates the signal, with strong suppression on larger scales, while PCA fails to account for the strong mode-mixing component on smaller scales. Hence, neither method adequately captures the foreground covariance at this level of contamination. Foreground avoidance with a 0.1 Mpc-1 buffer also fails to recover the signal for $k\leq 0.2$ Mpc-1. However, 1D PS results as shown in Figure  10 combining avoidance with a 0.1 Mpc-1 buffer along with GPR subtraction, though still inaccessible for $k<0.1$ Mpc-1, perform better on smaller scales compared to hybrid PCA ($n_{\text{fg}}=3$).
Figure 10: Spherically averaged 1D PS after residual foreground mitigation for 10% gain calibration error. Hybrid: The cyan markers indicate avoidance using a 0.1 Mpc-1 buffer. Blue and green markers show hybrid PCA and GPR results. The grey-shaded region represents the true HI signal while the brown, crossed-out regions represent the inaccessible $k$-modes due to avoidance with a 0.1 Mpc-1 buffer.
@@
As can be seen in Figure 11, the frequency-frequency covariance map of the foreground model retrieved after GPR exhibits smoothly varying features across the bandwidth, reminiscent of foregrounds, for a 1% gain calibration error, as opposed to comparatively less smooth features for a 10% error. For both contamination levels, incorporating additional mode-mixing kernels (e.g., MAT-3/2 and 5/2) into the GPR model further suppressed large-scale power on smaller scales, while improving the Bayesian evidence and recovering $l_{21}$ more accurately.
@@
Figure 11: Frequency-frequency covariance of the GPR-reconstructed residual foregrounds for 1% (left) and 10% (right) gain errors. Spectral smoothness degrades at higher gain errors, evident from diagonally auto-correlated structures in the 10% case.
@@ 5 Summary and Conclusion
In this study, we examined the impact of such gain errors on statistical detection of the 21-cm signal and assessed mitigation strategies based on residual foreground techniques. We used the 21cmE2E pipeline to simulate the mock observation data for SKA1-Low AA* configuration. We computed 2D and 1D PS from image cubes following three mitigation strategies: foreground subtraction (PCA, GPR), foreground avoidance, and a hybrid combination. Key findings for different gain error levels are summarised below:
  • Gain Error $\sim 0.1$%: Foreground avoidance about the horizon line enables successful recovery of the Hi signal across $k$-modes, with only $\sim$5% SNR loss. Subtraction or hybrid methods were unnecessary in this low-contamination regime.
  • Gain Error $\sim 1$%: GPR outperforms PCA on larger scales ($k\leq 0.1$ Mpc-1), but both methods fall short on smaller scales. Foreground avoidance with a 0.1 Mpc-1 buffer, or hybrid mitigation with a 0.05 Mpc-1 buffer post-subtraction, enables signal recovery across accessible $k$-modes, albeit with an average SNR loss of approximately 30%.
  • Gain Error $\sim 10$%: Subtraction (PCA or GPR) and avoidance (0.1 Mpc-1 buffer) alone fail to recover the signal. GPR especially overestimates small-scale power while suppressing large-scale modes. However, hybrid mitigation using GPR combined with avoidance successfully recovers the signal in the range $0.1\leq k\leq 0.5$ Mpc-1. Again, this recovery comes at an average 30% loss of SNR in the accessible $k$-modes.
Our results show that increasing gain errors contaminate the EoR window and erode the spectral smoothness of foregrounds, thereby limiting the effectiveness of standard subtraction-based methods. This raises a key question for future investigations: to what extent do foregrounds retain their smoothness when subjected to realistic instrumental effects and calibration systematics? Although the standard GPR technique [34] has the capability to model non-smooth foreground components, kernel selection remains a challenging task, especially in high-contamination scenarios [81]. Our findings suggest that a hybrid approach, combining model-based subtraction with strategic avoidance, can recover the signal Hi even under moderate to severe contamination. However, this comes at the cost of reduced power spectrum sensitivity and signal-to-noise ratio on cosmologically relevant scales. Therefore, careful inspection of the wedge boundary and the spectral windowing becomes essential [82].
Future work will include extending this framework to account for other observational systematics such as thermal noise, source position offsets, primary beam uncertainties, and bandpass calibration errors. We also plan to move toward Bayesian parameter estimation using emulator-based inference pipelines (e.g., [58]), and apply our methodology to real datasets, including upcoming Band-2 (120-250 MHz) uGMRT observations at $z=5.3,6.3,\,9.0\,\text{and}\,10.2$ (Sagar et al., in preparation).
@@ Acknowledgments
Software: We would like to acknowledge the usage of PYTHON programming language ( https://www.python.org/ ), numpy ([83]), scipy ([84]), astropy ([85]), emcee ([86]), and matplotlib ([87]) in this work.

### v2

@@ 3.3 Gaussian Process Regression (GPR)
where $\sigma^{2}$ is the variance, $\ell$ is the correlation lengthscale, and $\eta$ governs smoothness. Common choices include: RBF ($\eta\rightarrow\infty$), MATÉRN52 ($\eta=5/2$), MATÉRN32 ($\eta=3/2$), and exponential ($\eta=1/2$). For this work, we use RBF for sky foregrounds, Mat32/52 for mode-mixing, and exponential for the Hi signal. We note that although the standard GPR technique [34] has the capability to model non-smooth foreground components, kernel selection remains a challenging task, especially in high-contamination scenarios [82]. Hence, generic prior covariance functions like RBF or MAT32/52 may not always accurately represent the true covariances, which can lead to biases in the results in the low SNR regime [83]. Although the newer renditions of GPR like VAE-GPR [84, 80] use EoR simulations to train the kernels, this is beyond the scope of this work.
@@ 3.3.1 Application
0.1% 1% 10% 0.1% 1% 10%
$\sigma_{\text{sky}}/\sigma_{n}$ $\infty$ $\mathbb{U}(0.1,5)$ $\mathbb{U}(0.1,1)$ $\mathbb{U}(0.1,5)$ 1.02 1.03 0.48
$l_{\text{sky}}$ (MHz) $\mathbb{U}(5,50)$ $\mathbb{U}(5,50)$ $\mathbb{U}(1,25)$ 46.62 24.11 3.03
[…]
$l_{21}$ (MHz) $\Gamma(0.5,5)$ $\Gamma(0.5,5)$ $\Gamma(0.5,5)$ 0.51 26.39 11.53
Table 2: The GPR kernel priors and posteriors for 0.1%, 1% and 10% gain calibration error cases.
Figure 7: Posterior distributions of GPR hyperparameters after MCMC sampling for 0.1% gain calibration error.
@@ 4 Results
4.1 Case I: Gain Error of 0.1%
@@
In this case, we incorporate a gain calibration error of 0.1% and corrupt the simulated observed visibilities accordingly. The impact of a 0.1% gain calibration error on 1D PS is presented in Figure 5. In addition, the absolute fractional error of 2D PS, computed using Equation 2.7, is shown in the top panel of Figure 8. The 1D PS indicates that on larger scales ($k\leq 0.2$ Mpc-1), the residual power closely matches the target signal level, even without applying any mitigation. However, on smaller scales ($k>0.2$ Mpc-1), the residual power underestimates the true signal by nearly an order of magnitude.
[…]
4.2 Case II: Gain Error of 1%
@@
Figure 8: The 2D absolute fractional error maps $\Delta P_{i}(k_{\perp},k_{\parallel})$ for 0.1% (top row), 1% (middle row) and 10% (bottom row) gain calibration errors. Columns (from left to right) correspond to sky-model subtraction, PCA subtraction, and GPR subtraction. The dotted black lines and the shaded regions within indicate the wedge boundaries and the masked $k$-modes used for avoidance for that particular case. The vertical red dotted line marks the scale at which PSF begins to dominate. Bluer regions indicate better agreement with the true Hi signal.
@@
To further quantify the impact of calibration systematics, we introduce a gain calibration error of 1% and corrupt the simulated visibilities accordingly. The resulting cylindrical power spectrum is shown in the middle panel of Figure 4, while the corresponding spherically averaged power spectrum is presented in Figure 5. A comparison between the residual power spectrum, affected by a 1% gain calibration error, and the true Hi power spectrum reveals substantial discrepancies across all $k$ modes, with residual power consistently exceeding the true signal by a significant margin.
@@
The plots in the middle row in Figure 8 show the variation in 2D absolute fractional error maps for 1% calibration gain error, after PCA subtraction, and after GPR subtraction, respectively. In contrast to the 0.1% error case, one can see that the subtraction techniques effectively reduce the absolute fractional error by up to an order of magnitude in the low $k_{\perp}$ range ($\leq 0.1$Mpc-1). Although GPR performs marginally better for higher $k_{\perp}$ modes than PCA, the mode-mixing component dominates, and we do not observe a significant improvement with either method. The corresponding spherically averaged power spectrum after the subtraction techniques is presented in the left plot of Figure 9.
[…]
@@
Figure 9: Spherically averaged 1D PS after residual foreground mitigation for 0.1% (top-row), 1% (middle-row) and 10% (bottom-row) gain calibration error, respectively. Left (Subtraction only): Star markers denote PCA and GPR-subtracted residuals. The shaded region represents the true Hisignal. Right (Hybrid): The cyan markers show avoidance about the respective wedge boundaries (indicated by purple and yellow dotted vertical lines); green and blue markers represent hybrid (subtraction + avoidance) cases. The vertical red dotted line marks the scale at which PSF begins to dominate. Markers for GPR are offset by 5% along the $k$-axis for clarity.
4.3 Case III: Gain Error of 10%
@@
In this case, we corrupted the observed visibilities with a 10% gain calibration error to assess its impact on the recovered power spectrum. The effects are illustrated in the cylindrical power spectrum (bottom panel of Figure 4) and the spherically averaged power spectrum (Figure 5). We further examined the effectiveness of subtraction-based mitigation techniques in this high-error scenario. The plots in the bottom row in Figure 8 show the variation in the 2D absolute fractional error maps for 10% calibration gain error and after using subtraction methods. Applying GPR and PCA independently shows that subtraction alone is insufficient. GPR grossly overestimates the signal, with strong suppression on larger scales, while PCA fails to account for the strong mode-mixing component on smaller scales (left plot in the bottom-row of Figure 8). Hence, neither method adequately captures the foreground covariance at this level of contamination. Foreground avoidance with a 0.1 Mpc-1 buffer also fails to reduce the residual power to the signal levels for $k\leq 0.2$ Mpc-1. However, 1D PS results in the right plot bottom-row of Figure  9 shows that combining avoidance with a 0.1 Mpc-1 buffer along with GPR subtraction, though still inaccessible for $k<0.1$ Mpc-1, perform better on smaller scales compared to hybrid PCA ($n_{\text{fg}}=3$).
@@
As can be seen in Figure 10, the frequency-frequency covariance map of the foreground model retrieved after GPR exhibits smoothly varying features across the bandwidth, reminiscent of foregrounds, for a 1% gain calibration error, as opposed to comparatively less smooth features for a 10% error. For both contamination levels, incorporating additional mode-mixing kernels (e.g., MAT-3/2 and 5/2) into the GPR model further suppressed large-scale power on smaller scales, while improving the Bayesian evidence and recovering $l_{21}$ more accurately (see Appendix A).
@@
Figure 10: Frequency-frequency covariance of the GPR-reconstructed residual foregrounds for 1% (left) and 10% (right) gain errors without any mode-mixing components. Spectral smoothness degrades at higher gain errors, evident from diagonally auto-correlated structures in the 10% case.
@@ 5 Summary and Conclusion
In this study, we examined the impact of gain calibration errors on statistical detection of the 21-cm signal and assessed the efficacy of foreground mitigation strategies on corrupted synthetic observations. We used the 21cmE2E pipeline to simulate the mock observation data for SKA1-Low AA* configuration. We computed 2D and 1D PS in the range of 0.05 $\leq$ k $\leq$ 0.5 Mpc-1 from image cubes following three mitigation strategies: foreground subtraction (PCA, GPR), foreground avoidance, and a hybrid combination. It should be noted that we do not include galactic diffuse foregrounds, or frequency-dependent calibration gain errors in our simulations. Key findings for different gain error levels are summarised below:
[…]
Our results show that increasing gain errors contaminate the EoR window and erode the spectral smoothness of extragalactic foregrounds, thereby limiting the effectiveness of standard subtraction-based methods. This raises a key question for future investigations: to what extent do extragalactic foregrounds retain their smoothness when subjected to realistic instrumental effects and calibration systematics? Our findings suggest that a hybrid approach, combining model-based subtraction with strategic avoidance, can recover the signal Hi even under moderate to severe contamination. However, this comes at the cost of reduced power spectrum sensitivity and signal-to-noise ratio on cosmologically relevant scales. Therefore, careful inspection of the wedge boundary and the spectral windowing becomes essential [86].
Future work will extend this framework to account for additional observational systematics, including thermal noise, source position offsets, primary beam uncertainties, and bandpass calibration errors. We also plan to move toward Bayesian parameter estimation using emulator-based inference pipelines (e.g., [64]), and apply our methodology to real datasets, including upcoming Band-2 (120-250 MHz) uGMRT observations at $z=5.3,6.3,\,9.0\,\text{and}\,10.2$ (Sagar et al., in preparation).
@@ Acknowledgments
Software: We acknowledge the usage of PYTHON programming language ( https://www.python.org/ ), numpy ([87]), scipy ([88]), astropy ([89]), emcee ([90]), and matplotlib ([91]) in this work.
Appendix A GPR with mode-mixing
We present GPR results with a 1% calibration error after accounting for mode-mixing kernel effects when generating the foreground model. As one can see from the 1D plot, Figure 11, there is over-subtraction on small $k$-modes (Figure 15). The Figures 13 and 12 show the 2D PS and 1D PS of the GPR foreground models with (RBF + MAT52) and without (only RBF) a mode-mixing prior. The mode-mixing foreground model is picking up extra powers $k_{\perp}>0.1Mpc^{-1}$, as compared to the non-mode-mixing model.
Figure 11: Replication of Figure 9 but with mode-mixing effects included.
Figure 12: The 1D PS of foreground model generated by GPR with (green markers) and without (pink markers) considering a mode-mixing prior.
Figure 13: The 2D PS of foreground model generated by GPR with (right) and without (left) considering a mode-mixing prior.
We present posteriors after MCMC for 1% and 10% calibration gain errors in the 14. Comparing with the posteriors obtained in the mode-mixing plus smooth sky foreground case (15), we see that $l_{21}$ is constrained closer to its true value when we take mode-mixing effects into account.
Figure 14: Posterior distributions of GPR hyperparameters without mode-mixing kernels after for 1% (left) and 10%(right) gain calibration errors. Contours represent confidence levels of 11.8%, 39.3%, 67.5%, and 87.4%.
Figure 15: Posterior distributions of GPR hyperparameters, including mode-mixing kernel for 1% gain calibration error.